OpenStudy (anonymous):

The volume of a hexagonal prism is 10square root 3 in3. What is the height of the prism if the base area is 5square root 3 in3? A)50^square root 3 in B)2 in C)50 in D)2^square root 3 in

OpenStudy (anonymous):

V = H*B So H = V/B Now if we have this: \[H = \frac{ V }{ B } = \frac{ 10\sqrt 3 \in^3}{ 5 \sqrt 3 \in^2} = \frac{ 10 }{ 5 } = 2 inches\]

OpenStudy (anonymous):

Oops: Just take the \[\in = INCHES\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

No problem kiddo.